x^2 + 2x= 3/4=0
answer x= -3/2 or -1/2

need a full break down of how my book got this answer

Answer by Alan3354(69443)

x^2 + 2x= 3/4=0
answer x= -3/2 or -1/2
-----------
You have 2 equal signs. Try again.